Group Selection Page HELP Controls Group ITDM Vacuum WP78 FXE SPB HED SQS SXP Sample Environment PSPO XO TS Migrated2Zulip Archived2GitlabPages Archived2PDF MigratedByGroup
SCS Photon Commissioning Team XRO DOC WP74 MID XO_mig XO_Operation SQS Migrated2Zulip DET
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 23     Entry time: 13 Feb 2024, 06:38
Author: Giuseppe Mercurio 
Group: SCS 
Subject: Shift summary 

shift summary
Andreas, Giuseppe

  • continued pumpprobe measurements
  • around 5:30 we could not see any image at the MTE3 camera (after sample transfer), we tried to fix it but did not work
  • we tried to remove the samples from the LLC, the LLC is vented but the lift cannot be moved, needs to be repaired during the day.
ELOG V3.1.4-7c3fd00